Find the center and radius of the sphere

x2+6x+y2+18y+z2+6z=−50
Answer #1

Page No.: Date: _! Given 2? +6x + y² + 18y + 2? +62 = -50 x² + y2 + 2? +6X + 184 +67 + S0 = 0 - O_ the The Genera form of the